A 0.010 kg arrow is fired from a bow whose string exerts a force of 110 N on the arrow over a distance of 0.80 m. What is the SPEED of the arrow as it leaves the bow?
1
Expert's answer
2020-03-30T07:52:05-0400

The change of energy is equal to work done

"K_f-K_i=W=Fd""\frac{mv_f^2}{2}-0=Fd"

Hence, the final speed of the arrow

"v_f=\sqrt{\frac{2Fd}{m}}=\sqrt{\frac{2\times 110\times 0.80}{0.010}}=133\:\rm m/s."